A reaction of 7.20 mL of Iron(III) chloride solution (0.40 g/mL) and 8.95 g of K2C2O4.H2O (Potassium oxalate monohydrate) produced 6.22 g of K3Fe(C2O4)3.3H2O [Potassium trioxalatoferrate(III) trihydrate]

A) Calculate number moles of Iron(III) chloride?

B)Calculate number moles K2C2O4.H2O (Potassium oxalate monohydrate).

C)What is limiting reagent of the reaction (Consider stochiometry)?

D)Molar mass of Potassium trioxalatoferrate(III) trihydrate

E)Calculate the percent yield of the synthesis?
Answer #1

Balanced reaction given below
FeCl3 + 3K2C2O4·H2O ---> K3(Fe(C2O4)3)·3H2O + 3KCl

given

volume of FeCl3 = 7.2 ml
density = 0.4 g/ml
SO mass of FeCL3 present = density *volume= 7.2 ml*0.4 g/ml
= 2.88 g

molar mass of FeCL3 =1*55.84 g/mol + 3*35.45 g/mol= 162.2 g/mol
Number of moles present = mass/molar mass= 2.88g/162.2 g/mol =0.0177 mols

B)
given mass of K2C2O4.H2O = 8.95 g

molar mass of K2C2O4.H2O = [2*39.10 + 2*12.01+5*16 + 2*1.01} g/mol
= 184.23 g/mol

Number of mols of K2C2O4.H2Opresent = 8.95 g/184.23 g/mol =0.04858 mols

*****
C) 1 mole FeCl3 need 3 mol K2C2O4 H2O
So 0.0177 molFeCl3 need 3* 0.0177 mol K2C2O4 H2O = 0.0531
But moles ofK2C2O4 H2O =0.04858 mols
which is fewer .Hence K2C2O4 H2O act as limiting reactant.The reaction will stop once
all K2C2O4 H2O is consumed
limiting reactant =K2C2O4 H2O

******

D)
molar mass of K3(Fe(C2O4)3)·3H2O
=[3*39.10 +1*55.84+6*12.01+15*16+6*1.01 = 491.26 g/mol

****
limiting reactant =K2C2O4 H2O so it limits the reaction. It decide
amount of product formed
we have mole ratio K2C2O4 H2O :K3Fe(C2O4)3.3H2O = 3:1

SO moles of K3Fe(C2O4)3.3H2O produced = 1/3*0.04858 mols = 0.01619 mols
mass of K3Fe(C2O4)3.3H2O produced = number of moles of K3Fe(C2O4)3.3H2O *molar mass of
K3Fe(C2O4)3.3H2O = 0.01619 mols*491.26 g/mol =7.95 g

This is the theoretical yield ofK3Fe(C2O4)3.3H2O
actual yield is 6.22 g

Percent yield = actual yield/theoretical yield *100 = 6.22/7.95 *100 = 78.24%
***********
hope it is helpful